# An Empirical Analysis on the Use and Reporting of National Security Letters

### v3

@@ 4 Data Inconsistencies
4.1 Diverging Requests for Information Counts
The Office of the Director of National Intelligence (ODNI) started to publish an Annual Statistical Transparency Report (ASTR) in 2014 [34] including the number of issued NSLs and–similar to FISA [35]–NSL requests. The ASTR reports were a reaction to the first leaked documents by Snowden [40], ordered in June 2013 [16] to increase transparency.
The columns “FISA ROIs” and “ASTR ROIs” of Table 1 show the reported number of NSL requests. The “FISA ROIs” column is the sum of the reported NSL requests for information for US persons, non-US persons, and subscriber information under FISA in the previous three columns. Before 2015, the FISA numbers did not include non-US persons and subscriber information and can, therefore, not be compared to the ASTR numbers. However, we were unable to find any public statement (or observation) of the differing numbers in 2016, 2019, and 2021. It seems unlikely that there was a difference in the definition or counting of ROIs as they match exactly for the other years. While the numbers for 2016 and 2019 differ by less than 1%, there is an almost 19% deviation for 2021.
@@ 4.1 Diverging Requests for Information Counts
Table 1: The number of reported Request of Information (ROI) according to different sources. We sum the first three columns to derive the total number of ROIs reported under FISA and compare them with the ASTR counts reported by the ODNI.
4.2 NSL Metadata
@@
File numbers are part of the scarce metadata of NSL letters. Figure 5 plots the file numbers (y-axis) over the issuance date of the letter, for all NSLs that were published by companies after their gag order was lifted. It also plots the cumulative number of total NSL requests reported by ASTR, and the number of issued NSLs [^11] The x-axis labels mark the start of a year and we place the cumulative count at the end of a year. For example, the number of NSLs issued in 2014 is added to the cumulative count on Dec 31, 2014. Hence, it appears closer to the label of 2015. . The dashed red lines are the linear regression of the respective data. NSLs were issued since their authorization in 1978 and, therefore, the reporting that started in 2005 respectively 2013 has an unknown offset defined by the sum of previous NSL counts of the same type.121212 Early issues with the FBI tracking system add further uncertainty. The OIG reports that NSLs have been issued under the case file numbers of another division [37]. For ease of comparison with our hypothesis, we set the y-axis offset for these lines to the last previously known file number (for instance, for data starting in 2013, the last file number of an NSL from 2012 defines the y-axis offset).
@@
Figure 5: The file numbers of NSLs plotted over time show a monotonically increasing sequence of points. We compare them to the cumulative ROI and NSL counts published in the ASTR, after adjusting their y-axis offset to match the number of files in 2014. File numbers with red crosses have erroneous issue dates reported by companies.
@@
We note that the file numbers are increasing roughly monotonic where 22% of the numbers deviate and show slight decreases of 0.4% on average. It seems to be a reasonable hypothesis that NSL letters get assigned consecutive numbers by the FBI before their issuance (with some being issued faster than others afterward). The blue line of issued NSLs as reported by ASTR grows slower than the number of files. It is conceivable that some of the file numbers are assigned to NSLs that the FBI prepared to issue but withdrew before serving them to a company. [^13] A reason for more file numbers than issued letters would be if an investigation takes an unexpected turn that makes the NSL obsolete. For instance, when the FBI extracts the requested data directly from the laptop of the suspect with forensic methods, then there is no reason to serve the NSL and it might not be counted in the annual reports despite that it was already assigned a file number in the issuing process.
@@
Qualitatively, we observe that there appears to be some correlation between the cumulative number of ROIs and the number of files, delayed by a few months. The file numbers have two steeper increases, one from 2014–2015 and another from 2018–2019. The ROI counts show similar characteristics in 2015 and 2019 (recall, the cumulative counts are reported closer to the ticks in 2016 and 2020). These increases in the ROI counts are mainly caused by the spikes in the NSL requests for non-US persons, as Table 1 shows. However, we observe the possible anomaly that these steeper increases in file numbers are not reflected in the reported number of NSLs, despite that they appear to be present in the cumulative ROI counts. The expectation would have been that the file numbers follow trends in the number of issued NSLs and not the ROIs since a single NSL can contain multiple ROIs but should have only one file number. One possible explanation could be a difference in the counting methodology of ROIs for non-US persons. However, without internal insight into the assignment of file numbers, we are limited to observing these unexplainable inconsistencies, which negatively affects the confidence in the accuracy of the reported data.
5 Conclusion
Over the years, multiple efforts have been made to increase the transparency and congressional oversight on the use of National Security Letters (NSLs). Congress has amended the five NSL statutes several times, introducing mandatory reporting requirements, dozens of companies now publish NSL data (both individual requests and aggregated counts) while pushing for permission to publish yet finer-grained information [33]. However, despite these efforts, it is still challenging to assess how much transparency has really been provided as the data collection is challenging. No single data source provides a comprehensive view of how NSLs are used, and the manual effort entailed in collecting, extracting, normalizing, and analyzing disparate reports is an implicit obstacle to a more holistic analysis. This, in turn, results in uncertainty about what the public can and cannot understand about the use of NSLs, or the effectiveness of the amendments to NSL statutes.
We address these issues by gathering and publishing the first comprehensive collection of publicly reported data on NSLs, leveraging three sources: government statistics (such as reports by the FBI to Congress under FISA [35]), company transparency reports, and NSLs themselves, that were published after their nondisclosure order was lifted.
Our curated data set allows us to empirically analyze the use of NSL letters and compare trends across different sources. For example, from this data, we show a rough (albeit incomplete) agreement between corporate transparency reports and government data. Further, we use the small fraction of published NSL letters to infer the gag order time, which peaked in 2012 and has decreased since. Finally, we highlight that while the number of Requests for Information (ROIs) for US persons increased rapidly until 2010, it has since been superseded by the large but varying number of ROIs for non-US persons.
We advocate for publishing NSL statistics in a standardized and easily processable format (e.g., following the suggestions from [23]). This would simplify analyzing the NSL usage and trends that we showed in this paper to be insightful, and support Congress and the public in fulfilling their role to audit the government agencies. Furthermore, it would allow for earlier detection of internal and external inconsistencies, such as the discrepancy in FISA and ASTR reports, as well as the side-channel information from the file numbers that is inconsistent with the reported number of issued NSLs.
6 Acknowledgments
@@
The authors would like to thank the anonymous reviewers of CSLAW’24 for their insightful comments and suggestions. Furthermore, we thank Alisha Ukani and Stewart Grant for supporting us in interpreting our data with their data processing and visualization expertise.
Appendix A Data Sources
@@
A.1 List of Transparency Report Companies

### v4

@@ 5. Data Inconsistencies
5.1. Diverging Requests for Information Counts
The Office of the Director of National Intelligence (ODNI) started to publish an Annual Statistical Transparency Report (ASTR) in 2014 (astr, ) including the number of issued NSLs and—similar to FISA (fisa-reports, )—NSL requests. The ASTR reports were a reaction to the first leaked documents by Snowden (guardian-snowden, ), ordered in June 2013 (astr:2013, ) to increase transparency.
Table 1. The number of reported Request of Information (ROI) according to different sources. We sum the first three columns to derive the total number of ROIs reported under FISA and compare them with the ASTR counts reported by the ODNI. Numbers corrected after we reported the inconsistencies are shown in parentheses.
@@ 5.1. Diverging Requests for Information Counts
The columns “FISA ROIs” and “ASTR ROIs” of Table 1 show the reported number of NSL requests. The “FISA ROIs” column is the sum of the reported NSL ROIs for US persons, non-US persons, and subscriber information under FISA in the previous three columns. Before 2015, the FISA numbers did not include non-US persons and subscriber information and can, therefore, not be compared to the ASTR numbers.
There were no public statements (or observations) of the differing numbers in 2019 and 2021, despite that the number of ASTR ROIs is over 23% higher than the count of FISA ROIs. After reporting these inconsistencies to the ODNI, the National Security Division issued a correction of the ROIs for subscriber information from 2019, which should have been 19,061 instead of 19,601 (fisa-correction-2020, ). Moreover, we were informed that the numbers of the ASTR report for 2021 will be corrected when the ASTR report for 2024 is issued in Spring 2025. The correct number would have been 31,825, as reported by FISA.
5.2. NSL Metadata
@@
File numbers are part of the scarce metadata of NSL letters. Figure 5 plots the file numbers (y-axis) over the issuance date of the letter, for all NSLs that were published by companies after their gag order was lifted. It also plots the cumulative number of NSL requests reported by ASTR, and the number of issued NSLs. [^11] The x-axis labels mark the start of a year and we place the cumulative count at the end of a year. For example, the number of NSLs issued in 2015 is added to the cumulative count on Dec 31, 2015. Hence, it appears closer to the 2016 label. The dashed red lines are the linear regression of the respective data. NSLs were issued since their authorization in 1978 and, therefore, the reporting that started in 2005 and 2013 respectively have an unknown offset defined by the sum of previous NSL counts of the same type.121212 Early issues with the FBI tracking system add further uncertainty. The OIG reports that NSLs have been issued under the case file numbers of another division (ig-report-2014, ). For ease of comparison with our hypothesis, we set the y-axis offset for these lines to the last previously known file number (for instance, for data starting in 2013, the last file number of an NSL from 2012 defines the y-axis offset).
@@
Figure 5. The file numbers of NSLs plotted over time show a monotonically increasing sequence of points. We compare them to the cumulative ROI and NSL counts published in the ASTR, after adjusting their y-axis offset to match the number of files in 2014. File numbers with red crosses have erroneous issue dates reported by companies.
[…]
[Plot showing the file numbers respectively the cumulative count of the number of issued NSLs on the y axis and time on the x axis.] The file numbers of NSLs plotted over time show a monotonically increasing sequence of points. We compare them to the cumulative ROI and NSL counts published in the ASTR, after adjusting their y-axis offset to match the number of files in 2014. Some file numbers are outliers and marked with red crosses. They had erroneous issue dates reported by companies.
@@
We note that the file numbers are increasing roughly monotonic where 22% of the numbers deviate and show slight decreases of 0.4% on average. It seems to be a reasonable hypothesis that NSL letters are assigned consecutive numbers by the FBI before their issuance (with some being issued faster than others afterward). The blue line of issued NSLs as reported by ASTR grows slower than the number of files. It may be the case that some of the file numbers are assigned to NSLs that the FBI prepared to issue but withdrew before serving them to a company.
@@
Qualitatively, we observe that there appears to be some correlation between the cumulative number of ROIs and the number of files, delayed by a few months. The file numbers have two steeper increases, one from 2014–2015 and another from 2018–2019. The ROI counts show similar characteristics in 2015 and 2019 (recall, the cumulative counts are reported closer to the ticks in 2016 and 2020). These increases in the ROI counts are mainly caused by the spikes in the NSL requests for non-US persons, as Table 1 shows. However, we observe the possible anomaly that these steeper increases in file numbers are not reflected in the reported number of NSLs, despite appearing to be present in the cumulative ROI counts. It is expected that file numbers follow trends in the number of issued NSLs and not ROIs, since a single NSL can contain multiple ROIs but should have only one file number. One possible explanation could be a difference in the counting methodology of ROIs for non-US persons. The ODNI did not comment on this inconsistency. Without internal insight into the assignment of file numbers, we are limited to observing these unexplained inconsistencies, which lowers our confidence in the accuracy of the reported data.
6. Current Challenges and Potential Paths Forward
In this section, we start by discussing the current operational challenges towards providing better transparency. We then describe potential paths forward to improve upon the current state of affairs.
6.1. Understanding Current Challenges
We conducted two informal interviews with the current and former Chief, ODNI Civil Liberties, Privacy, and Transparency Office to validate our findings and understand the process and challenges of government data curation. These interviews were specific to this situation and the ODNI component that we engaged with. We do not know the extent to which the sentiments expressed to us fully generalize to other Federal agencies in the national security sphere. We summarize our understanding of this ODNI perspective on NSL reporting below.
An important challenge for the national security community is that publishing data about intelligence collection must be evaluated for its potential risk to security. As documented in ”Principles of Intelligence Transparency for the Intelligence Community” (odni:ic-principles, ) published by the ODNI, there is a requirement to balance between transparency and security to avoid accidental disclosure of sensitive information. Among the concerns expressed to us is metadata leakage—that a document may implicitly embed sensitive information in its digital metadata. Thus, as seen in our data, many NSLs are published as PDF scans, specifically to exclude a potential metadata information channel. Unfortunately, this also significantly hinders the parsing of the information that was intentionally released in the NSL report. Another concern is the unauthorized disclosure of classified information via the content itself (reflected in the redaction in some of the PDFs). Last but not least, publishing too much or too specific data might lead to accidental information leaks (e.g., that correlating or aggregating might allow a third-party to learn information that they should not be able to), which is a serious concern for the intelligence community.
Another practical challenge is the lack of a unified system and methodology for counting NSLs. On the one hand, definitions of what should be counted are sometimes vaguely defined in the statute, which can lead to ambiguity in counting even within a particular organization. On the other hand, different agencies reporting on the same data may have different methodologies for counting. Moreover, the current system for tracking NSLs is not set up to support counting.
A third challenge explained to us was the lack of resources for such government processes. It is already a labor-intensive process to collect the numbers and minimize the risk of accidental information leaks, not to mention conducting additional analysis to ensure correctness. Further exacerbating this issue, it was mentioned that different agencies sometimes need to report data over different time periods, which appears to limit the ability to correlate or normalize processes.
Finally, as a meta-challenge, we as independent researchers have limited ability to evaluate the choices or tradeoffs made here because the underlying security concerns are inherently opaque and difficult to judge without access to classified information.
In summary, ODNI believes that providing better transparency and data remains a challenging task, as it can require a delicate balancing between transparency, security, resource allocation, and other operational challenges. Moreover, identifying whether existing transparency mechanisms are achieving their policy goals and whether substantive expansion might be achievable without security impact, may require a perspective only available within the national security community. [^13] We wonder if an organization such as the Privacy and Civil Liberties Oversight Board (PCLOB) might not offer an appropriate vehicle for such an evaluation, as its charter allows to investigate such questions in this environment.
6.2. Potential Paths Forward
Given these limitations, and based on our insights from studying NSL data collection and its challenges, we make the following recommendations for the publication of future transparency data.
First, it is feasible to significantly increase the usability of existing published data, and hence the ability of the public to interpret and review published data, without jeopardizing the security goals of the intelligence community. We suggest publishing data in text files, as comma-separated values (the “CSV” file type). This document type has no metadata, but is machine-readable and, hence, easy to process. Intelligence.gov already maintains a public dataset, and it can be considered to expand this to other agencies. Indeed, this is how mandated wiretap reporting is structured141414www.uscourts.gov/data-news/reports/statistical-reports/wiretap-reports—a dataset with similar concerns about inadvertent disclosure. Second, we recommend the Congress consider consolidating data collection efforts. Currently, the DOJ and ODNI both collect and report NSL data. This not only duplicates counting efforts, but different reporting periods and ambiguous definitions—e.g., about scope and categorization of NSLs—may introduce discrepancies in the reported data. A single reporting agency could better normalize reporting and reduce challenges in interpretation. Third, NSL data reported under FISA has little context, which is essential for interpretation. For instance, our analysis showed that the number of requests for information fluctuate over the years. For instance, the 39,214 ROIs reported by ASTR in 2021 may sound plausible in isolation, but knowing that the previous year only had 24,225 may be helpful to spot anomalies. (And in this specific case, could have helped to identify that the reported number was indeed several thousand ROIs too high.) Without the historical data, longitudinal trends cannot be observed and the reported data cannot be put into perspective. If possible, commentary on anomalies over the years would help interpreting unexpected peaks.
We acknowledge that these changes may be challenging to achieve for existing processes. However, we hope our empirical analysis of the effectiveness of existing NSL transparency measures will inform future debates of the Congress on transparency regulation and reporting mandates. These discussions can be complemented by research on alternative data reporting requirements for optimal auditability as well as the government’s knowledge of what data is sensitive to national security.
7. Discussion and Conclusion
Transparency is an aspirational tool for discouraging the government from abusing its power. A common way to balance transparency with secrecy is through reporting sensitive data in aggregate forms. In this way, the public can review the broad actions of the government and flag significant violations without revealing the details of individual activities. However, the success of this approach largely relies on two assumptions being true—that the data is provided in a well-documented and usable fashion, and that the public (or portions thereof) can and is regularly auditing the data to ensure compliance.
In this work, we have explored the extent to which these two assumptions hold for NSL usage in practice. By compiling a broad array of data published by both the public and private sectors, we are able to holistically evaluate the current state of data transparency of NSLs. We further highlight that while data has been released to provide transparency to the public, the usability of the data leaves something to be desired. This lack of usability in turns hinders any public effort to audit the published data. Indeed, as part of the process, we have uncovered data discrepancies that were corrected after reporting.
Our findings suggest that both assumptions may be too strong in practice. Indeed, our work is not the only one to make such observations. Other researchers have identified concerns about the trustworthiness and completeness of the wiretap reports long mandated by 18 U.S.C § 2519 (see Gidari (web:wiretap-data, ) and Varner and Ng (news:missing-warrants, )). These concerns are echoed by the findings of an internal investigation conducted by the Federal Judicial Center, which found the government’s wiretap reports to be inaccurate and incomplete (fjc:wiretap-report, ). The underlying causes were multifaceted, but included a combination of procedure challenges, operational complexities, ignorance, and imperfect incentives. Thus, while transparency may have the potential to encourage compliance and enable public oversight, this promise is clearly predicated on effective execution—which may be undermined by a range of practical and organizational factors.
Acknowledgments
@@
We thank our anonymous reviewers and shepherd James Grimmelmann for their valuable comments and suggestions. We thank Alisha Ukani and Stewart Grant for supporting us in interpreting our data with their data processing and visualization expertise. Additionally, we are grateful to our interviewees at ODNI for their time and sharing their insights. Finally, we thank Peter Swire and Sunoo Park for their feedback on the draft.
Appendix A Additional Background for NSLs
In this section, we provide additional background information for NSLs.
A.1. Statutory Provisions for NSLs
At the time of writing, five statutory provisions grant government agencies legal authority to issue NSLs. Namely, the Right to Financial Privacy Act (RFPA) [4], (Title II of) the Electronic Communications Privacy Act (ECPA) [8], the National Security Act (NSA) [1], the Fair Credit Reporting Act (FCRA) [4], and the USA PATRIOT Act [16]. We introduce each statute in chronological order.
A.1.1. The Right to Financial Privacy Act of 1978
The RFPA, codified at 12 U.S.C. § 3401 et seq. [4], is the first statutory provision for NSLs. It defines a special procedure (codified at 12 U.S.C. § 3414 [3]) for government agencies to request the production and disclosure of financial records for the purposes of conducting intelligence activities. However, this legislation did not mandate access to such records. In practice, financial institutions declined requests from the FBI due to conflicts with State privacy or banking laws that prohibited the production of such records [6, 34]. This led to the amendment of the RFPA in 1986.
A.1.2. The Amendment to RFPA in 1986
In October 1986, as part of the Intelligence Authorization Act for Fiscal Year 1987 [10], Congress passed an act that amended RFPA to explicitly grant the FBI access to financial records provisioned in 12 U.S.C. § 3414 [7]. This amendment also introduced a mandate that required a semiannual report on the number of requests made to certain congressional bodies (the select committee of the House and the Senate) and a nonclosure requirement without a specific end date [10].
A.1.3. The Electronic Communications Privacy Act of 1986
In parallel with the amendment to RFPA, Congress passed the second statutory provision for NSLs as part of the Electronic Communications Privacy Act (ECPA) of 1986 [11]. The Title II of the ECPA, which is called the Stored Communications Act (SCA) and codified at 18 U.S.C. § 2701 et seq. [8], gave the FBI access to business records of wire or electronic communication service providers for counterintelligence purposes (18 U.S.C. § 2709 [9]). Furthermore, this act defined similar non-disclosure and reporting requirements to those in Section A.1.2.
A.1.4. The Amendment to NSA in 1994
First introduced in 1947, the National Security Act [1] was amended in 1994 to include the third statutory provision for NSLs (as part of the Intelligence Authorization Act for Fiscal Year 1995 [13]). This amendment introduced a procedure (codified at 50 U.S.C § 3162 [12]) for any authorized agency to request a broad array of business records from various organizations for investigating potential document leaks from government employees. Compared to the previous two statutory provisions, this one permitted more agencies to request a wider range of documents albeit only under specific circumstances. Additionally, unlike the first two NSL statues, it did not have a reporting mandate.
A.1.5. The Amendment to FCRA in 1996
Initially enacted in 1970, the Fair Credit Reporting Act (FCRA) [4] was amended in 1996 to incorporate the fourth statutory provision for NSLs (as part of the Intelligence Authorization Act for Fiscal Year 1996 [14]). This amendment included a procedure (codified at 15 U.S.C. § 1681u [2]) for the FBI to access credit agency records. Like those defined in RFPA and ECPA, it came with a reporting mandate for certain congressional bodies on a semiannual basis and allowed the FBI to demand nondisclosure.
A.1.6. The PATRIOT Act of 2001
The PATRIOT Act of 2001 [16] was enacted as a response to the terrorist attacks on September 11 [28]. It made substantial amendments to three of the four existing NSL statutes (RFPA, ECPA, and FCRA) and added a fifth one. The amendments in this act extended the scope of NSLs to allow the FBI to request information for people who are not themselves a foreign power or its agent but are relevant to an investigation concerning international terrorism or foreign spying. In addition to making NSLs more widely applicable, the amendments also simplified the approval process by authorizing Special Agents in Charge to issue NSLs in addition to officials in the FBI headquarters [34].
The fifth NSL statute was introduced by amending FCRA. It introduced a procedure (codified at 15 U.S.C. § 1681v [2]) for government agencies to access consumer reports from credit reporting agencies to conduct the investigations of intelligence or terrorist activities. Compared to other statutes, this fifth one came with a nondisclosure requirement but not a reporting mandate (such an obligation was later added as part of a subsequent amendment).
A.2. Subsequent Amendments
A.2.1. RFPA amendment in 2003
A.2.2. PATRIOT Act amendments
In 2006, the 109th Congress amended the USA PATRIOT Act with two statues, in part as a response to earlier judicial reactions to the USA PATRIOT Act [45]. Specifically, two court cases Doe v. Gonzales [58] and Doe v. Ashcroft [59]151515 The plaintiff’s name “John Doe” is used in both cases because the gag order prohibited the disclosure of the plaintiff’s identity. raised First and Fourth Amendment issues with NSLs connected to the non-disclosure provisions and the absence of judicial oversight [34]. In the case of Doe v. Ashcroft, the district court held that NSLs violated the Fourth Amendment because they authorized “coercive searches effectively immune from any judicial process” [61]. Moreover, the court held that the nondisclosure provisions unconstitutionally restricted free speech as they can prohibited disclosure without providing judicial means to challenge a ban or achieve eventual relief [61, 60].
As a reaction to both cases, and while Doe v. Ashcroft was on appeal at the Second Circuit, the 109th Congress amended the USA PATRIOT Act with the USA PATRIOT Improvement and Reauthorization Act [20] and the USA PATRIOT Act Reauthorization Amendments Act [19]. Collectively, besides adding penalties for non-compliance with NSLs or their nondisclosure requirements [18, 17], these two statutes introduced judicial review on the use of NSLs and weakened the nondisclosure requirements. Specifically, the amendments require that the Foreign Intelligence Surveillance Court (FISC) approves NSL requests and adds control mechanisms and a process to ease nondisclosure provisions [18, 17]. Furthermore, the changes mandated classified semiannual reports about the use of NSLs to Congress and nonclassified annual statistics reports [22]. The latter reports are one of the main data sources for this paper. Regarding Doe v. Ashcroft, the Second Circuit ruled in 2008 that the FBI needs to certify that disclosure of an NSL would lead to statutorily enumerated harms to justify non-disclosure provisions [44]. Last, the amendments call for an audit on use of the NSLs from the Office of the Inspector General (referred to as the OGI reports).
A.2.3. Office of the Inspector General (OIG) reports
As part of the PATRIOT Act amendments (P.L. 109-177 § 119 [21]), the Office of Inspector General (OIG) of the Department of Justice was required to audit and report on the use of NSLs. On March 9, 2007 [50], soon after the FBI started reporting data on NSLs, the OIG released its first report. Notably, the report suggested that the internal processes by which the FBI tracks NSLs were flawed. For instance, NSLs that are issued under different statutes are tracked independently which might lead to counting the same person multiple times. Similar issues were mentioned in follow-up reports, noting that the same person might be counted as multiple people due to inconsistent spelling of their name. In response to the OIG reports, the FBI took steps to correct the identified deficiencies and introduced a new internal database for tracking NSLs. While they tried to retroactively correct information about NSLs for 2006 and 2007 manually, they stated that this information might still be inaccurate [63].
A.2.4. USA FREEDOM Act
In 2015, Congress enacted the USA FREEDOM Act [27], which revised the nondisclosure requirement, introduced an additional reporting requirement, and expressly limited the use of NSL to specifically identified information. Notably, the additional reporting requirement and restrictions on the use of NSLs are direct responses to NSA’s bulk metadata collection practice [51].
Appendix B Data Sources
@@
B.1. List of Transparency Report Companies